Lilian9667:

2x-3=2x1-5 A.) -2 B.) 0 C.) 5/3 D.) no solution

supie:

The `1` is a variable ?

Lilian9667:

what

supie:

2x-3=2x1-5 ^ Is that supposed to be a variable?

supie:

i mean exponent* i get those mixed up e.e

Lilian9667:

Ooops 1 wasn't supposed to be there hehe

supie:

Oh, so it's supposed to be \(2x-3=2x-5\)?

Lilian9667:

yah lol

supie:

Lol ok So you know the first step?

Lilian9667:

This what I did 2x-3=2x-5 -6 -6 =4

Lilian9667:

know if that's right

supie:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @Lilian9667 This what I did 2x-3=2x-5 -6 -6 =4 \(\color{#0cbb34}{\text{End of Quote}}\) I'm confused on how you got that ._.

Lilian9667:

Yah me too um lol Hey i tried

supie:

I think the first step is to subtract \(2x\) from both sides of the equation \(2x−3=2x−5\) You think you can do that?

Lilian9667:

Oh I meant to put -10 on the right

supie:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @Lilian9667 Oh I meant to put -10 on the right \(\color{#0cbb34}{\text{End of Quote}}\) Uhh yea...no

Lilian9667:

I guess

supie:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @supie I think the first step is to subtract \(2x\) from both sides of the equation \(2x−3=2x−5\) You think you can do that? \(\color{#0cbb34}{\text{End of Quote}}\)

Lilian9667:

um

supie:

Ok well \(2x−3=2x−5\) then subtract \(2x\) \(2x−3−2x=2x−5−2x\) \(=?\)

Lilian9667:

ohhh

supie:

Yeah So that would that =?

supie:

ur stuck still?

Lilian9667:

yah

Lilian9667:

I'm trying to solve it on papper

supie:

Ok lmk when ur done and it's `paper*` btw

Lilian9667:

=6

supie:

How would that =6 There is two equations \(2x−3−2x=2x−5−2x\) Has to equal \(?=?\)

supie:

Not just one number youre subtracting 2x from both sides so youll have two

Lilian9667:

-2

supie:

I just told you it will have two numbers ?=?

Lilian9667:

oh well NEXT THIS IS CONFUSING

supie:

Ok \(2x−3−2x=2x−5−2x\) = \(−3=−5\)

supie:

Then the next step would be to do \(+3\) to both sides of the equation The equation \(−3=−5\)

supie:

So \(−3+3=−5+3\) equals what? It's a little bit like the last one

Lilian9667:

+6 +8

Lilian9667:

I think?

supie:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @Lilian9667 I think? \(\color{#0cbb34}{\text{End of Quote}}\) Nope it's wrong The it would actually be \(0=−2\)

supie:

You get it now?

supie:

So is there a solution or no?

Lilian9667:

I just said -2. lol. Yeah?

supie:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @Lilian9667 I just said -2. lol. Yeah? \(\color{#0cbb34}{\text{End of Quote}}\) uhh no you said "+6 +8"

Lilian9667:

nooo seee

Lilian9667:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @Lilian9667 -2 \(\color{#0cbb34}{\text{End of Quote}}\) You see that

Lilian9667:

I said that

supie:

oh...yeah but that was for the other one but ig it's still right you just skipped a step but the real correct answer is \(0=−2\) not just -2 though but ig its still right

Lilian9667:

Oh ok

supie:

ight so what is your answer from your answer options?

supie:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @supie ight so what is your answer from your answer options? \(\color{#0cbb34}{\text{End of Quote}}\) A.) -2 B.) 0 C.) 5/3 D.) no solution

Lilian9667:

need Layla to teach me to step by step on Linear equations

Lilian9667:

A.) -2

supie:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @Lilian9667 A.) -2 \(\color{#0cbb34}{\text{End of Quote}}\) Actually I'm not sure what is the answer they want because \(0=-2\) would be no solution but they can also want the answer –2

Lilian9667:

Ok :)